if the length of a rectangle is 12 units and its width is 9 units, what is the length of its diagonal?
Answer by jim_thompson5910(35256)

Use the pythagorean theorem

a^2 + b^2 = c^2

12^2 + 9^2 = c^2

144 + 81 = c^2

225 = c^2

c^2 = 225

c = sqrt(225)

c = 15

The length of the diagonal is 15 units